0.2g of sample of H2O2 required 10ml of 1N KMnO4 solution in a titration in presence of dil. H2SO4. What is the percentage of purity of H2O2?

Grade:11

1 Answers

ABHISHEK JAIN AskiitiansExpert-IITD
13 years ago

dear Abhingya,

equivalent weight of H2O2 =  eq. weight of KMnO4

W/E = N × V

W/17 = 1 × 10/1000

W = .17

Hence % of purity = .17/.2 × 100

                          = 85 %

 

with regards,

abhishek jain